### v1

# Is fixed-node diffusion quantum Monte Carlo reproducible?

Abstract
Fixed-node diffusion quantum Monte Carlo (FN-DMC) is a widely-trusted many-body method for solving the Schrödinger equation, known for its reliable predictions of material and molecular properties. Furthermore, its excellent scalability with system complexity and near-perfect utilization of computational power makes FN-DMC ideally positioned to leverage new advances in computing to address increasingly complex scientific problems. Even though the method is widely used as a computational gold standard, reproducibility across the numerous FN-DMC code implementations has yet to be demonstrated. This difficulty stems from the diverse array of DMC algorithms and trial wave functions, compounded by the method’s inherent stochastic nature. This study represents a community-wide effort to address the titular question, affirming that: Yes, FN-DMC is reproducible (when handled with care). Using the water-methane dimer as the canonical test case, we compare results from eleven different FN-DMC codes and show that the approximations to treat the non-locality of pseudopotentials are the primary source of the discrepancies between them. In particular, we demonstrate that, for the same choice of determinantal component in the trial wave function, reliable and reproducible predictions can be achieved by employing the T-move (TM), the determinant locality approximation (DLA), or the determinant T-move (DTM) schemes, while the older locality approximation (LA) leads to considerable variability in results. This work lays the foundation to establish accurate and reproducible FN-DMC estimates for all future studies across applications in materials science, physics, chemistry, and biology.
1 Introduction
The credibility of a scientific result hinges on its reproducibility; independent and equivalent experiments should lead to the same conclusion. Achieving reproducibility is, however, not easy. There are several historical examples from both social and natural sciences[1, 2, 3, 4] that have served to illustrate its challenges, and substantial ongoing effort is dedicated to addressing this so-called “reproducibility crisis” [5, 6]. The problem of reproducibility is particularly pertinent within computational experiments in the hard sciences, where different computational codes should ideally lead to the same prediction. Nonetheless, reproducibility can be compromised by small algorithmic differences, undocumented approximations, and undetected bugs in the simulation software or its dependencies (numerical libraries, compilers etc.). Determining the source of discrepancies can be difficult, e.g., due to restricted source code availability[7, 2, 8, 9].
Here, we consider reproducibility in the context of the many-electron Schrödinger equation[10], fundamental to the quantum mechanical description of matter, and its countless applications to physics, chemistry, biology, engineering, and materials science. In this context, the topic of reproducibility has been recently addressed[11, 12] in two seminal papers for density functional theory (DFT) – the work-horse of materials science. However, despite its widespread success, DFT often falls short of providing the necessary quantitative, and sometimes qualitative, description of key complex systems. Fortunately, advances in hardware, algorithms, and fundamental theories are paving the way for the routine application of methods beyond the accuracy of DFT. The scope of these methods has recently broadened significantly beyond simple benchmarks, towards an extensive description of molecules, surfaces and condensed phases[13, 14, 15, 16, 17, 18] that can include complex dynamics facilitated by machine learning potentials[19, 20, 21, 22, 23, 24, 25, 26].
Fixed-node diffusion Monte Carlo[27, 28, 29] (FN-DMC) is an accurate state-of-the-art computational approach for solving the Schrödinger equation for a variety of systems, including molecules, solids, and surfaces. This method obtains the electronic ground-state by performing an imaginary-time evolution from a starting trial wave function . Within the Born-Oppenheimer approximation, consists of the real space positions of all the electrons. Typically, is the product of an antisymmetric function (e.g., a Slater determinant or a sum of Slater determinants[30]) and a symmetric, positive function, called the Jastrow factor[31]. The Jastrow factor is explicitly dependent on electron-electron and electron-nucleus distances, and able to directly capture a significant fraction of the electronic correlation.
The FN-DMC projection is achieved with an ensemble of electron configurations, known as walkers, which evolve according to the imaginary-time Green function[32], yielding a drift-diffusion process over discrete imaginary time steps, , to stochastically sample the ground-state wave function; the stochastic uncertainty is then inversely proportional to the square root of the number of samples. The main approximation in FN-DMC is that the fixed-node wave function is constrained to have the same nodal surface as , in order to avoid the so-called fermion sign problem. [33] This introduces a variational error in the computed ground state energy. For single-reference systems, this error is typically small even for simple single determinant trial wave functions built from DFT. FN-DMC exhibits almost perfect efficiency on modern supercomputers[34, 35, 36] and a cubic scaling per Monte Carlo step with system size[37], making it often the only computationally affordable method beyond DFT for treating large condensed phase systems with more than 100 atoms. Over time, numerous algorithmic improvements have enhanced the accuracy, efficiency, and stability of FN-DMC. These advances have enabled the successful application of FN-DMC to a wide array of problems across the natural sciences, including the calculation of the energies of condensed phases and large molecules[18, 14, 15, 38, 39, 40], the binding of molecules on surfaces[17, 41, 42, 43, 44, 45, 46], phase diagrams[47, 48, 49, 50, 20, 51, 52, 53], reaction barrier heights[54, 55, 56, 57, 58], spin-polarized uniform electron gas[59], two-dimensional electron liquid[60], lithium systems[61], electronic and optical properties of delafossites[62], defect formation energies[63, 64], calculation of energy derivatives[65, 66, 67], radical stabilization energies[68], excited states[69, 70, 71, 72, 73, 74, 75, 76, 77, 78], training of quantum machine learning models[79], electron-positron interactions[80], polymorphism[81, 82, 83], electronic band gaps[84], Landau-level mixing in quantum dots[85], localization in quantum dots and quantum wires[86, 87, 88, 89], nearly exact density functional quantities[90, 91] and more. Recent progress in the use of neural networks as trial wave functions for FN-DMC[92, 93, 94] has served to boost its accuracy and potential future uptake even further.
There exist several codes that have been under active development for decades, each of which makes somewhat different algorithmic and implementation choices, such as different Jastrow factors. This diversity raises open questions on the reproducibility of FN-DMC. If FN-DMC is to be widely accepted as a highly accurate reference method, it is important that consistent results can be obtained from these different FN-DMC codes. With this goal in mind, the present letter represents a collaborative effort among the users and developers of eleven distinct FN-DMC codes, to rigorously assess the reproducibility of FN-DMC.
A key obstacle to the reproducibility of FN-DMC comes from the use of non-local pseudopotentials (NLPPs), which increase the efficiency of the method for systems with heavy atoms. [95, 96, 97] NLPPs allow one to solve the Schrödinger equation solely for the valence electrons, by substituting the full local nuclear potential with a smooth non-local potential near the nuclei. In general, NLPPs hinder reproducibility in electronic structure methods, as NLPPs constructed in different ways can lead to somewhat different predictions. This issue is even more challenging in FN-DMC because non-local pseudopotential operators create an additional sign problem in the projector beyond the one that is always present for fermionic calculations. To avoid this sign problem, these operators must be “localized”,[98] or at least partially localized,[99] on a wave-function. A natural choice is to localize them on the trial wave-function , introducing a dependence on both the determinantal and the Jastrow components of the wave function. Since the Jastrow factor is different in the different codes and its parameters are stochastically optimized, yielding possible noise and reproducibility issues, some authors choose to localize only on the determinantal component. [100, 96, 101, 102, 103] This removes the dependence on the Jastrow factor at the cost of losing the desirable property that the treatment of the pseudopotential is exact in the limit of exact . To summarize, there are currently four localization schemes: the locality approximation (LA)[104, 98], the T-move (TM) approximation[99, 105, 106], the determinant locality approximation (DLA)[100, 96, 101, 102, 103], and the determinant T-move (DTM) approximation[103]. These four schemes (LA, TM, DLA and DTM) result in somewhat different projected wave-functions and therefore different total energies of physical systems.
In this work, we systematically compare the four localization algorithms across eleven FN-DMC codes (named alphabetically): Amolqc, CASINO[34], CHAMP-EU[107], CHAMP-US[108], CMQMC, PyQMC[109], QMC=Chem[110, 111], QMCPACK[36, 112], QMeCha[113], QWalk[114], and TurboRVB[35, 115]. Different forms of Jastrow factor are necessarily tested as part of this evaluation. We specifically consider the case of the interaction energy between methane and water molecules. The methane-water dimer is characterized by weak binding (), which makes it a highly stringent target and thus a great test for reproducibility. In addition, it is a prototype of more complex systems such as methane clathrates, important for gas storage and transportation. We show that consensus across all eleven codes can be made when utilizing the TM, DLA and DTM approximations, particularly following careful control of the discretized time step.
2 Results and Discussion
First, we compute the interaction energy of the methane-water dimer using the eleven codes for the four different localization schemes (where available). The interaction energy of the methane-water dimer,
is defined as the difference between the energy of the complex, , minus the sum of the energies of the isolated water and methane monomers (see the Methods section for details on the geometries and the DMC simulation set-up). All the interaction energies are extrapolated to the zero time-step limit according to the procedure described in the supporting information (SI) and in Ref. [116].
We note that two results are reported for the TurboRVB code, namely TurboRVB (DMC) and TurboRVB (LRDMC). TurboRVB (DMC) refers to the standard FN-DMC algorithm with time step discretization and available with the T-move scheme. However, production simulations of FN-projection in TurboRVB are typically performed with the lattice regularized DMC (LRDMC)[117, 105], which is an alternative approach to DMC. In particular, LRDMC is based on a lattice regularization of the many-electron Hamiltonian over a spatial mesh, and the ground state is projected out via the Green function Monte Carlo method[118, 119, 120]. The zero mesh-size limit of the LRDMC prediction is equivalent to the zero time-step limit of DMC, and is therefore also included in this work. We also note that the T-move approximation itself comes in four different versions as briefly discussed in the SI but, when presenting the TM results, we will not distinguish between them because they differ only at finite time step, while we report here the extrapolated values at zero time step, where they are equivalent.
The computed methane-water interaction energies are shown in Fig. 1. We plot the FN-DMC interaction energy computed with each code with a colored circle. In addition, the average among the interaction energies computed with different codes is reported with a gray dashed line, and its statistical error with a shaded gray region. The average value and its statistical error are computed as the mean value and the standard deviation of the probability distribution reported in Eq. 2, discussed later on in the manuscript. We compare the prediction of FN-DMC to the value computed by coupled cluster theory with single, double, triple, and perturbative quadruple excitations [CCSDT(Q)], which is expected be highly accurate for weak intermolecular interactions[121] (details of the calculation are reported in Sec. S3 of the SI). Despite using only a single determinant in the trial wave functions and a DFT nodal surface for simplicity, broadly speaking, the FN-DMC is in excellent agreement with CCSDT(Q) (black dashed line). However, a strikingly large spread of predictions across different codes is obtained when using the LA. In contrast, the TM, DLA, and DTM methods show a much narrower spread of the interaction energies.
The data reported in Fig. 1 allows us to estimate a probability distribution of the interaction energy for each analyzed method. In particular, we write the DMC energy estimated with the code and the method (LA, TM, DLA, DTM) as , and its statistical error bar as . Following the central limit theorem, we expect each DMC estimate to be distributed according to a normal distribution, with mean and standard deviation . Since we do not know and , we use here the current estimates and and define the probability distribution of the energy for the method as:
where is the number of codes for which the localization method is evaluated. The mean and the variance of the energy for the distribution are respectively:
and
In particular, the variance takes into account both the statistical error bar of each FN-DMC evaluation () and its deviation from the mean value ().
The probability distributions are plotted in Fig. 2. When the LA is employed, the probability distribution is spread across a large energy range of , with a standard deviation of meV. The agreement across different codes significantly improves with the TM, DLA and DTM schemes, with the probability distributions showing a quite localized peak (standard deviation of ca. meV or less) centered on meV, meV and meV respectively. The DTM scheme gives the narrowest distribution, centered on meV, with a standard deviation of ca. 1 meV, but since only four out of the eleven codes implemented DTM this is not of much significance. Overall, the analysis of the probability distributions showcases that algorithms more sophisticated than LA need to be employed to guarantee reproducibility among different FN-DMC codes.
A key factor in DMC is the convergence with respect to the simulation time step. The projection is only accurate for sufficiently small time step, requiring calculations at various time steps to be performed and extrapolated to the limit . The required time step depends on both the system being studied and the accuracy of the trial wave function. For this reason, we also analyze the dependence of the probability distribution on the simulation time step and report it in Fig. 3. In particular, we consider the case of the DLA, for which we have computed the interaction energy with several codes at multiple time steps (). We notice that, for a large time step , the DLA energy predictions are spread across a large energy range of over . Decreasing the time step leads to a significant reduction in the distribution’s variance. At the time step of , the probability distribution becomes very narrow, indicating agreement among different codes. We highlight here that the converged time step is system-dependent, and the time step behavior is highly sensitive to different codes and approximations, as shown in the SI. Therefore, an analysis of the convergence with respect to the simulation’s time step is important to achieve a converged and reproducible FN-DMC energy, and a fair comparison across different packages.
To conclude, this work establishes FN-DMC as a robust reference method by demonstrating reproducibility across 11 popular codes with varying implementations. This work marks a significant milestone for FN-DMC and the greater electronic structure theory community, giving confidence in the multidisciplinary array of applications of FN-DMC. Crucially, we show that reliable and reproducible FN-DMC predictions can be achieved when using the TM, DLA, or DTM approximation schemes, rather than the LA scheme, for pseudopotential simulations while ensuring convergence with respect to the time step. Notably, we are able to achieve agreement to within a standard deviation of meV on the interaction energy of the methane-water complex — which is approximately two hundred thousand times smaller than the total energy of the dimer. This achievement lays the groundwork for more extensive studies in the future, including calculations on the physical and chemical properties in the condensed phase.
3 Methods
The interaction energy of the methane-water dimer is computed by subtracting the isolated molecule energies from the methane-water complex, as defined in Eq. 1. The geometry of the dimer (shown in the SI) was obtained from Ref. [122]. The geometries of the monomers are the same as in the dimer. In this study, in order to try to achieve consistent results, all eleven codes were required to use the same correlation consistent effective core potential (ccECPs)[123, 124] and the corresponding triple-zeta basis set (ccECP-ccpVTZ), as well as a Slater-Jastrow wave function with a single Slater determinant whose orbitals are obtained from DFT calculations using the Perdew-Zunger parametrization[125] of the local-density approximation. For the methane-water dimer, this was sufficient to obtain accurate results. Some of the codes exchanged wave function data via the TREXIO library. [126] Every code implements a slightly different parametrization of the Jastrow factor, but all codes include in the Jastrow factor an electron-electron (e-e), an electron-nucleus (e-n), and an electron-electron-nucleus (e-e-n) term. The variational parameters of the Jastrow factor have been optimized by minimizing either the variational energy or the variance, according to the recommended scheme within each code. The time steps employed in each simulation are in the range 0.001 to 0.1 a.u. The final estimates reported in Fig. 1 were extrapolated to the limit using the procedure described in the SI. Further details specific to each code, the schemes used to deal with the localization error, the time step extrapolation, and the tests on the size consistency error are reported in the SI.
Acknowledgements
M.B. acknowledges the computational resources from the HPC facilities of the University of Luxembourg [127] (see hpc.uni.lu). M.D. acknowledges financial support from the European Union under the LERCO project number CZ.10.03.01/00/22_003/0000003, via the Operational Programme Just Transition, and computational resources from the IT4Innovations National Supercomputing Center (e-INFRA CZ, ID: 90140). M.C. acknowledges access to French computational resources at the CEA-TGCC center through the GENCI allocation number A0150906493. J.T.K., P.R.C.K., Y.L. and L.M. were supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, as part of the Computational Materials Sciences Program and Center for Predictive Simulation of Functional Materials. Part of the work of L.M. has been supported also by U.S. National Science Foundation grant DMR-2316007 and employed resources at NERSC at early stages for this project. M.C., E.S., R.S., and C.F. acknowledge partial support by the European Centre of Excellence in Exascale Computing TREX — Targeting Real Chemical Accuracy at the Exascale. This project has received funding in part from the European Union’s Horizon 2020 — Research and Innovation program — under grant agreement no. 952165. E.S., R.S., and C.F performed the calculations on the Dutch national supercomputer Snellius with the support of SURF Cooperative. K.N. acknowledges financial support from the JSPS Overseas Research Fellowships and from MEXT Leading Initiative for Excellent Young Researchers (Grant No. JPMXS0320220025) and computational resources from the Numerical Materials Simulator at National Institute for Materials Science (NIMS). L.K.W and W.W. were supported by U.S. National Science Foundation via Award No. 1931258. A.Z. and D.A. acknowledges support from the European Union under the Next generation EU (projects 20222FXZ33 and P2022MC742). Y.S.A., D.A. and A.Z. acknowledge support from Leverhulme grant no. RPG-2020-038. A.M. and B.X.S. acknowledge support from the European Union under the “n-AQUA” European Research Council project (Grant No. 101071937). The portion of the work done by T.A.A. and C.J.U. received initial support under AFOSR (Grant No. FA9550-18-1-0095) and was completed under the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. Department of Energy Office of Science and the National Nuclear Security Administration."
This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. Calculations were also performed using the Cambridge Service for Data Driven Discovery (CSD3) operated by the University of Cambridge Research Computing Service (www.csd3.cam.ac.uk), provided by Dell EMC and Intel using Tier-2 funding from the Engineering and Physical Sciences Research Council (capital grant EP/T022159/1 and EP/P020259/1), and DiRAC funding from the Science and Technology Facilities Council (www.dirac.ac.uk). This work also used the ARCHER UK National Supercomputing Service (https://www.archer2.ac.uk), the United Kingdom Car Parrinello (UKCP) consortium (EP/ F036884/1).